# Semantic Communication and Control Co-Design for Multi-Objective Correlated Dynamics

### v1

@@ III-B Logic Semantic Koopman Model
In this phase, we assume that all plants, except plant $i$, follow predefined control rules. Plant $j$ (where $\𝑖j\in\mathcal{I}\backslash i$) operates under control rule $\varphi_{j}$ and exhibits different dynamics from plant $i$. Using the encoder, decoder, and DSK model from plant $i$, plant $j$ derives an LSK model (represented by matrices $\mathbf{L}^{A}{\varphi_{j}}$ and $\mathbf{L}^{B}{\varphi_{j}}$), which captures the system dynamics under $\varphi_{j}$. The optimization problem for the LSK model is given as
$\underset{\mathbf{L}^{A}_{\varphi_{j}},\mathbf{L}^{B}_{\varphi_{j}}}{\arg\min}\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}\left\lVert\mathbf{x}_{i,k}-\Psi^{-1}\left(\mathbf{L}^{A}_{\varphi_{j}}\mathbf{D}_{A}\mathbf{z}_{i,k}+\mathbf{L}^{B}_{\varphi_{j}}\mathbf{D}_{B}\mathbf{u}_{i,k}\right)\right\rVert^{2}_{2}.$ (16)
where $\mathbf{L}^{A}_{\varphi_{j}}$ and $\mathbf{L}^{B}_{\varphi_{j}}$ represent the Koopman matrices parameterized by the control rules $\varphi_{j}$. Due to challenges in obtaining an analytical solution, deep learning is employed. A new loss function, logic loss ($\mathcal{L}_{4}$), ensures the state representations follow the control rule $\varphi_{j}$ and is given by
@@
Once the LSK model is fully trained, plant $j$ integrates it with the DSK model from plant $i$ to construct its Koopman model and perform remote control operation.
@@ IV Simulation Results
This section will validate the proposed logical Koopman AE approach on inverted cart-pole systems over wireless channels. Each system is defined by a four-dimensional state vector representing the horizontal position and velocity of the cart and the vertical angle and angular velocity of the pendulum, and control is applied through a horizontal force. Here, we adopt the following simulation parameter settings unless specified otherwise: the pendulum mass is set to $\{1,3,4,5,6\}\,\mathrm{Kg}$, the cart mass is established at $\{5,15,20,25,30\}\,\mathrm{Kg}$, and the pendulum length is maintained at $\{0.2,0.6,0.8,1.0,1.2\}\,\mathrm{m}$. We simulate five systems with varying parameters, governed by logical control rules, over a temporal span of $[150,251]$. These logical control rules are defined as
@@
for $i\in\{1,2,3,4,5\}$. The parameter $\beta_{i}$, ranging across $\{0,2.0,3.0,4.0,5.0\}$, delineates the value within the state space for each respective system. The training data is sampled at a rate of $\tau_{0}=100\,\mathrm{ms}$ over the time interval of $\left[0,251\right]$ and system dynamics are computed using the Runge-Kutta method with a step size of $0.1$. The logical Koopman AE model is trained using the Adam optimizer with a batch size of $251$, a learning rate of $2\times 10^{-3}$, and early stopping, and experiments are repeated five times with different seeds.
The encoder has a fully connected layer of $32$ neurons, activated by a ReLU function, followed by an output layer of $d\in\{2,4\}$ neurons with sigmoid activation. The decoder mirrors the encoder, while the DSK model includes two fully connected layers, and the LSK model adds two more layers after the DSK matrices are applied. The model is trained with weight hyperparameters $c_{1}=0.1$, $c_{4}=10^{-6}$, and $c_{2}=c_{3}=0.75$ for (14), and $c^{\prime}_{1}=0.1$, $c^{\prime}_{5}=10^{-6}$, $c^{\prime}_{2}=c^{\prime}_{3}=0.75$, and $c^{\prime}_{4}=0.1$ for (18). Wireless performance is evaluated based on SNR values $\gamma_{i,k}\in\{5,15\}\,\mathrm{dB}$ with a transmission bandwidth ($B$) of $20\;\mathrm{MHz}$, a path loss exponent ($\eta$) of $3.0$, a transmission power ($P$) of $0.1\mathrm{Watt}$, and the distance between communication pairs ($D$) of $50\mathrm{m}$.
The performance is assessed by sample complexity (number of communication samples required to train the model) and prediction accuracy using normalized root mean squared error (NRMSE), computed as
$\mathcal{N}_{i,K_{p}}=\frac{\sqrt{\frac{1}{K_{p}}\sum_{k=K_{s}+1}^{K_{s}+K_{p}}\|\tilde{\mathbf{x}}_{i,k}-\mathbf{x}_{i,k}\|^{2}_{2}}}{\|\max(\mathbf{x})-\min(\mathbf{x})\|_{2}}\times 100,$ (20)
where $K_{p}$ is the prediction horizon at the test. System stability is measured by a score function that assigns a score of $1$ when the cart’s position is within $0.2$ units of the target location and the pendulum’s angle remains within $0.05$ radians of the upright position and 00 otherwise We compare the logical Koopman AE model with two baselines:
@@
Impact of Representation Dimensions.  Fig. 3 presents the state prediction performance and sample complexity of the proposed logical Koopman AE model with two baseline models, using an SNR of $15$ dB, one-step prediction depth, and varying state representation dimensions. The logical Koopman AE model performs similarly to Baseline 1 and significantly outperforms Baseline 2 across all systems. It achieves an NRMSE of about $1\%$ as shown in Fig. 3(a), with the reference system (System 1) reaching $0.2\%$ NRMSE with a four-dimensional representation. Lower-dimensional representations slightly increase the NRMSE to $1.5\%$ as shown in Fig. 3(b), while Baseline 2 shows a much higher $9\%$ error.
@@
The proposed model’s success comes from its ability to encode control rules in the latent space via the LSK model, pre-training the encoder and decoder for $50$ epochs. The DSK model captures the semantic correlations leveraging reference dynamics, requiring only $10\%$ of the data to encode control rules compared to reference dynamics. By combining DSK and LSK, the model effectively linearizes non-linear dynamics, enabling accurate state predictions with low communication overhead. Note that the encoder and decoder are fine-tuned on systems governed by control rules, enriching the latent space to capture dynamic variations.
Baseline 2, trained only on the reference system, struggles with prediction accuracy, especially with fewer communication samples, due to its inability to handle systems governed by different control rules. Fig. 3(a) shows that high-dimensional representations further improve state prediction performance compared to low-dimensional representations by capturing the most critical eigenfunctions of the system dynamics and modeling the control rules in the latent space. Baseline 1 also benefits from a larger representation dimension, improving Koopman matrix modeling to predict future states at the cost of increased communication overhead. Reducing from four to two-dimensional representations can cut communication samples in half while maintaining reasonable prediction accuracy, hence it is necessary to evaluate control performance under different representation dimensions.
Control Performance.  Fig. 4 shows the control performance of the proposed model with two baselines, measured by average score values under different representation dimensions and an SNR of $15$ dB. The proposed model closely matches the reference values across all systems, as does Baseline 1, which maintains strong prediction accuracy (around $0.5\%$ NRMSE for four-dimensional and $1\%$ for two-dimensional representations). However, Baseline 2 shows lower state prediction accuracy, impacting its control performance. Despite slight reductions in prediction accuracy with two-dimensional representations shown in Fig. 4(b), the control performance remains nearly unchanged, suggesting that lower-dimensional representations can reduce communication without sacrificing control performance. Average score values tend to decrease for systems further from the reference, as they take longer to reach the target state, leading to slight control performance degradation.
@@
Figure 4: Average score values for each system of proposed and baselines with $\text{SNR}=15$ dB for (a) $d=4$ and (b) $d=2$.
Figure 5: State prediction and communication samples for each system of proposed and baselines for (a) $\text{SNR}=15$ dB and (b) $\text{SNR}=5$ dB.
Impact of SNR.  Fig. 5 shows the state prediction performance and sample complexity of the proposed model with two baselines across different SNR values using two-dimensional representations and one-step prediction depth. The results indicate that, for both models, state prediction accuracy decreases slightly at lower SNR values compared to higher ones. Both models experience a slight decrease in accuracy at lower SNR values, primarily due to increased packet loss at $5$dB, which reduces the training data available. Interestingly, Baseline 1 maintains similar or better performance for systems $4$ and $5$ at low SNR, likely due to the smaller, more balanced dataset preventing overfitting. In contrast, the proposed model’s accuracy declines more significantly at low SNR, likely due to error accumulation in its DSK and LSK components.
@@ V Conclusion
This letter introduces a novel Logical Koopman AE framework designed to manage multi-objective distinct control systems, departing from the previous work that focused on controlling multi-objective identical control systems. The framework integrates two key models: the DSK model, which captures the reference system’s semantic structure and linearizes state evolution in the latent space, and the LSK model, which encodes control rules for systems with different control rules and dynamics. Results show that the framework reduces communication overhead for controlling multi-objective distinct dynamics and improves state prediction accuracy without compromising control performance.

### v2

@@ III-B Logic Semantic Koopman Model
[…]
$\underset{\mathbf{L}^{A}_{\varphi_{j}},\mathbf{L}^{B}_{\varphi_{j}}}{\arg\min}\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}\left\lVert\mathbf{x}_{i,k+k^{\prime}}-\Psi^{-1}\left(\tilde{\mathbf{z}}_{i,k+1}\right)\right\rVert^{2}_{2}.$ (17)
with the future predicted state representation of the plant $j$ at time $k+1$ given as
$\tilde{\mathbf{z}}_{i,k+1}=\mathbf{L}^{A}_{\varphi_{j}}\mathbf{D}_{A}\mathbf{z}_{i,k}+\mathbf{L}^{B}_{\varphi_{j}}\mathbf{D}_{B}\mathbf{u}_{i,k},$ (18)
where $\mathbf{L}^{A}_{\varphi_{j}}$ and $\mathbf{L}^{B}_{\varphi_{j}}$ represent the transformation matrices parameterized by the control rules $\varphi_{j}$. Due to the challenges in obtaining an analytical solution, deep learning is employed. A new loss function, logic loss ($\mathcal{L}_{4}$), ensures that state representations follow the control rule $\varphi_{j}$ and is given by
@@
Once the LSK model is fully trained, plant $j$ integrates it with the DSK model to construct its Koopman model and perform remote control operations.
@@ IV Simulation Results
This section validates the semantic logical Koopman AE on inverted cart-pole systems over wireless channels. Each system is defined by a four-dimensional state vector representing the horizontal position and velocity of the cart and the vertical angle and angular velocity of the pendulum, and control is applied through a horizontal force. Here, we adopt the following simulation parameter settings unless specified otherwise: the pendulum mass is set to $\{1,3,4,5,6\}\,\mathrm{Kg}$, the cart mass is established at $\{5,15,20,25,30\}\,\mathrm{Kg}$, and the pendulum length is maintained at $\{0.2,0.6,0.8,1.0,1.2\}\,\mathrm{m}$. We conduct simulations on five distinct systems, each characterized by varying parameters that change their underlying dynamics, and governed by logical control rules, over a temporal interval $[150,251]$. These logical control rules are defined as
@@
for $i\in\{1,2,3,4,5\}$. The parameter $\beta_{i}$, ranging across $\{0,2.0,3.0,4.0,5.0\}$, specifies the value within the state space for each respective system. The training data is sampled at a rate of $\tau_{0}=100\,\mathrm{ms}$ over a trajectory length ($K$) of $\left[0,251\right]$, and the system dynamics is computed using the Runge-Kutta method with a step size of $0.1$. The semantic logical Koopman AE is trained using the Adam optimizer with a batch size of $256$, a learning rate of $2\times 10^{-3}$, and early stopping, and the experiments are repeated five times with different seeds.
The encoder consists of a fully connected layer with $h_{1}=32$ neurons, activated by a ReLU function, followed by an output layer with $d\in\{2,4\}$ neurons and sigmoid activation. The decoder mirrors the encoder, while the DSK model includes two fully connected layers, and the LSK model adds two more layers after the DSK matrices are applied. The model is trained with weight hyperparameters $c_{1}=0.1$, $c_{4}=10^{-6}$, and $c_{2}=c_{3}=0.75$ for (15), and $c^{\prime}_{1}=0.1$, $c^{\prime}_{5}=10^{-6}$, $c^{\prime}_{2}=c^{\prime}_{3}=0.75$, and $c^{\prime}_{4}=0.1$ for (20). Wireless performance is evaluated based on SNR values $\gamma_{i,k}\in\{5,15\}\,\mathrm{dB}$ with a transmission bandwidth ($B$) of $20\;\mathrm{MHz}$, a path loss exponent ($\eta$) of $3.0$, a transmission power ($P$) of $0.1\,\mathrm{Watt}$, and the distance between communication pairs ($D$) of $50\,\mathrm{m}$.
The performance is evaluated by sample complexity (number of communication samples required to train the model) and prediction performance using normalized root mean square error (NRMSE), calculated as
$\mathcal{N}_{i,K_{p}}=\frac{\sqrt{\frac{1}{K_{p}}\sum_{k=K_{s}+1}^{K_{s}+K_{p}}\|\tilde{\mathbf{x}}_{i,k}-\mathbf{x}_{i,k}\|^{2}_{2}}}{\|\max(\mathbf{x})-\min(\mathbf{x})\|_{2}}\times 100\%,$ (22)
where $K_{p}$ is the prediction horizon at the test. The control performance is measured by a scoring function that assigns a score of $1$ when the cart’s position is within $0.2$ units of the target location and the pendulum’s angle remains within $0.05$ radians of the upright position, and 0 otherwise. We compare the semantic logical Koopman AE with two baselines:
@@
Impact of Representation Dimensions.  Fig. 3 illustrates the state prediction performance and sample complexity of the proposed semantic logical Koopman AE in comparison with two baselines, evaluated under an SNR of $15\,\mathrm{dB}$, a one-step prediction depth, and varying state representation dimensions. The semantic logical Koopman AE achieves performance comparable to Baseline 1 and significantly outperforms Baseline 2 across all systems. As shown in Fig. 3(a), the proposed model achieves an NRMSE of about $1\%$, with the reference system (system 1) reaching $0.2\%$ with a four-dimensional representation. Reducing the representation to two dimensions slightly increases the NRMSE to $1.5\%$ as shown in Fig. 3(b), while baseline 2 exhibits a substantially higher error of $9\%$.
@@
The improved performance of the proposed model is attributed to its ability to capture semantic structure through the DSK model trained on the reference system, combined with the LSK model, which encodes system-specific control rules in the latent space. The encoder and decoder are pre-trained for $50$ epochs using only $10\%$ of the data to train the LSK model compared to the DSK model. By jointly leveraging the DSK and LSK models, the proposed approach effectively linearizes non-linear system dynamics, enabling accurate state prediction with significantly reduced communication overhead. Note that the encoder and decoder are fine-tuned on each target system to adapt to system-specific control rules, enriching the latent space to capture dynamic variability.
Although the proposed model reduces the communication cost by transmitting and predicting low-dimensional representations, it introduces a modest computational cost that stems from encoding logical control rules in the latent space. Its training complexity combines two parts: the DSK and LSK models. Specifically, the DSK training scales as $O(N.\mathcal{E}.(p.h_{1}+h_{1}.p+d^{2}))$, where $\mathcal{E}$ is the number of epochs during DSK training and $N$ is the number of training samples, while the LSK training scales as $O(0.1N.\mathcal{E}_{l}.d^{2})$, where $\mathcal{E}_{l}$ is the number of epochs during LSK training. Hence, the overall training complexity is $O(N.\mathcal{E}.(p.h_{1}+h_{1}.p+d^{2})+0.1N.\mathcal{E}_{l}.d^{2})$. During the encoding of logical control, the operational time complexity of encoding control rules over trajectories of length $K$ scales as $O(0.1N.\mathcal{E}_{l}.K.|\varphi|)$, where $|\varphi|$ is the number of operators in the logical control formula. Despite this computational overhead, the proposed model is more efficient than Baseline $2$, which requires training a Koopman AE separately for each system, incurring per-system cost comparable to the DSK model training. In contrast, the proposed model trains the DSK model once and, for each new system, incurs only the small additional cost of training the lightweight LSK model. This trade-off enables efficient scalability across heterogeneous systems while maintaining low communication and computation requirements.
As shown in Fig. 3(a), high-dimensional representations enable the proposed model to capture dominant Koopman eigenfunctions, improving state prediction performance. In contrast, reducing the representation dimension from four to two reduces the number of transmitted samples by approximately half while maintaining acceptable prediction accuracy.
Control Performance.  Fig. 4 presents the control performance of the proposed model and two baselines, evaluated in terms of the average score values under varying representation dimensions and SNR of $15\,\mathrm{dB}$. The proposed model achieves control performance closely aligned with the reference values across all systems, similar to baseline 1, which maintains high prediction accuracy with approximately $0.5\%$ NRMSE for four-dimensional and $1\%$ for two-dimensional representations. In contrast, baseline 2 exhibits lower prediction accuracy, which negatively impacts its control performance.
Although reducing the representation dimensions to two leads to a slight decline in prediction accuracy, the control performance remains nearly unaffected, as shown in Fig. 4(b). This result suggests that lower-dimensional representations can substantially reduce communication overhead without compromising control performance. Furthermore, average score values decrease for systems that deviate more significantly from the reference system, as these systems require more time to reach the desired state, leading to control performance degradation.
@@
Figure 4: Average score values for each system of the proposed and baselines with $\text{SNR}=15$ dB for (a) $d=4$ and (b) $d=2$.
Figure 5: State prediction and communication samples for each system of the proposed and baselines for (a) $\text{SNR}=15$ dB and (b) $\text{SNR}=5$ dB.
Impact of SNR.  Fig. 5 illustrates the state prediction performance and sample complexity of the proposed model compared to two baselines for different SNR values, two-dimensional representations, and one-step prediction depth. The results indicate that the prediction accuracy slightly decreases with lower SNR values for all models, mainly due to increased packet loss at $5\,\mathrm{dB}$, which reduces the training data. Interestingly, baseline 1 maintains comparable improved performance for systems $4$ and $5$ under low SNR values. This behavior is likely attributed to a smaller and more balanced dataset that mitigates overfitting. In contrast, the proposed model exhibits a more noticeable decline in prediction accuracy at low SNR values, which can be attributed to error propagation within the DSK and LSK models.
@@ V Conclusion
In this work, we proposed a novel semantic logical Koopman AE designed to control multi-objective distinct systems under limited communication and computational resources. We employed two key models: the DSK model to capture the shared semantic structure in a linear form within the latent space, and the LSK model to encode system-specific control rules within the latent space. Simulation results validate that the proposed approach’s ability to scale across multi-objective distinct systems without compromising control performance.